# Time-Dependent Radiation Quality Factor $\langle Q\rangle$ of Galactic Cosmic Rays in Deep Space and Shielding Environments: Modeling and Measurements

### v1

@@ 1 Introduction
    The Cosmic Ray Telescope for the Effects of Radiation (CRaTER, Spence et al., 2010; Mazur et al., 2011) on board the Lunar Reconnaissance Orbiter (LRO, Chin et al., 2007; Vondrak et al., 2010), launched in 2009 and orbiting the Moon, has offered invaluable data on the lunar radiation environment, including the LET spectrum and absorbed dose rate, for over 15 years (e.g., Schwadron et al., 2012; Zeitlin et al., 2013a, 2019).
@@ 3.3 Validations against CRaTER and TGO Measurements
We then examine the temporal variations of the absorbed dose rate, dose equivalent rate, and averaged radiation quality factor, and present their evolutions alongside the SSN and $\phi$ values in Figure 6. Panel (a) shows the monthly SSN and the corresponding solar modulation potential adopted in the BON20 model, together illustrating the evolution of solar activity and associated modulation of GCRs. Panel (b) compares the modeled and measured absorbed dose rates $D$ for CRaTER and TGO/Liulin-MO. For the CRaTER D1 and D2 pair, the modeled absorbed dose rate (blue) generally aligns with the measured trend (purple), and the vertical bars in the modeled curve represent uncertainties arising from lunar albedo contributions (8.6–19.9%; Schwadron et al., 2012; Zaman et al., 2020), with the central value corresponding to the median albedo contribution of 14.25%. Both the modeled and measured $D$ values decrease during solar maximum and increase toward solar minimum, consistent with the expected GCR modulation. A generally closer match is obtained for CRaTER before 2019, while noticeable discrepancies appear afterward, likely because the BON20 calibration precedes this period and 2019–2021 appears to be at an unusually quiet solar minimum. For TGO/Liulin-MO, the modeled (orange) and measured (brown) absorbed dose rates at Mars’s orbit show overall aligned temporal variations. In the modeled results, in addition to the correction factors for ACR presence and the radial gradients of GCRs and ACRs in A, albedo particle contributions are accounted for using the fractions presented in Section 3.3 of Liu et al. (2023), which vary in the range of 4–8% depending on the solar modulation at an altitude of 400 km around Mars. As the modeled albedo particle contributions in Liu et al. (2023) provide specific values, the associated uncertainties are not shown here. However, additional uncertainties may arise from the variability of the model inputs (e.g., SSN) and from the regolith composition and physical properties of the Martian surface (e.g., Röstel et al., 2020), which influence the generation and transport of albedo particles and is beyond the scope of this study. The measured data are converted to dose rates in water with a factor of 1.333, as derived in Appendix B of Schwadron et al. (2012). Both modeled and measured values for TGO/Liulin-MO are systematically higher than those for CRaTER, due to (1) enhanced contributions from secondary particles under the thicker shielding environment (see also, e.g., Dobynde et al., 2019; Xu et al., 2022; Zhang et al., 2022; Liu et al., 2023), and (2) radial gradient of the GCR and ACR fluxes between Earth and Mars, as detailed in A.
@@ 4.1 Radiation Dose Contribution from Individual GCR Elements
In summary, the modulation of $\langle Q\rangle_{Z}$ with solar activity depends strongly on elemental mass and charge. In order to elucidate the underlying physics of the contrasting behavior mentioned above, we analyze how $\langle Q\rangle)i$ evolves with solar modulation for light and heavy species in the following.
@@ Acknowledgments
The authors acknowledge the support by the National Natural Science Foundation of China (Grant Nos. 42521007, 42188101 and 42474221). The authors also express their gratitude to the groups developing the BON20 GCR model and the GEANT4 toolkit, the World Data Center SILSO, Royal Observatory of Belgium, Brussels, and the teams of ACE/CRIS, AMS-02, LRO/CRaTER and TGO/Liulin-MO instruments for providing valuable data involved in this study.
@@ Appendix A Calculations of ACR Contributions, and GCR and ACR Radial Gradients
Note that BON20 only incorporates GCRs, whereas ${D}_{\text{BON20}}^{*}(t)$ in Equation (3) includes both ACR and GCR contributed dose:
@@
and the dose equivalent $H$ is corrected in the same manner. For convenience, we compute the time-averaged correction factors for the GCR and ACR contributions in Equation (13) and apply these averaged factors to absorbed dose and dose equivalent estimates in deep space or under different shielding conditions:
$C_{r,\,{\rm GCR+ACR}}=\left\langle\frac{D_{\rm GCR}(r,t)}{D_{\rm GCR}(r=1\,\mathrm{AU},t)}\right\rangle_{t}+\left\langle\frac{D_{\rm ACR}(r,t)}{D_{\rm ACR}(r=1\,\mathrm{AU},t)}\right\rangle_{t}.$ (14)
At $r=1.5$ AU ($\Delta r=0.5$ AU), the correction yields a 1.75% increase in both $D$ and $H$ due to GCR gradients, and ($9.5\pm 0.06$)% and ($8.2\pm 1.3$)% increases in $D$ and $H$, respectively, due to ACR contributions. Interestingly, at 1 AU, Equation (14) gives ($5.2\pm 0.06$)% and ($5.7\pm 0.7$)% increases in $D$ and $H$, respectively, consistent with the 6.3% estimate from Equation (5), confirming that including only He, N, and O provides a reasonable approximation of total ACR effects. Therefore, we adopt correction factors of 11.0% for $D$ and 9.7% for $H$ for long-term deep-space and shielded dose rates at Mars, although some deviations may occur for different shielding thicknesses.

### v2

# Time-Dependent Radiation Quality Factor $\langle Q\rangle$ of Galactic Cosmic Rays in Deep Space and Shielded Environments: Modeling and Measurements

@@ 1 Introduction
    The Cosmic Ray Telescope for the Effects of Radiation (CRaTER, Spence et al., 2010; Mazur et al., 2011) on board the Lunar Reconnaissance Orbiter (LRO, Chin et al., 2007; Vondrak et al., 2010), launched in 2009 and orbiting the Moon, has offered invaluable data on the lunar radiation environment, including the LET spectrum and absorbed dose rate, for over 15 years (e.g., Schwadron et al., 2012; Zeitlin et al., 2013a, 2019; Looper et al., 2020).
@@ 3.3 Validations against CRaTER and TGO Measurements
We then examine the temporal variations of the absorbed dose rate, dose equivalent rate, and averaged radiation quality factor, and present their evolutions alongside the SSN and $\phi$ values in Figure 6. Panel (a) shows the monthly SSN and the corresponding solar modulation potential adopted in the BON20 model, together illustrating the evolution of solar activity and associated modulation of GCRs. Panel (b) compares the modeled and measured absorbed dose rates $D$ for CRaTER and TGO/Liulin-MO. For the CRaTER D1 and D2 pair, the modeled absorbed dose rate (blue) generally aligns with the measured trend (purple), and the vertical bars in the modeled curve represent uncertainties arising from lunar albedo contributions (8.6–19.9%; Schwadron et al., 2012; Zaman et al., 2020), with the central value corresponding to the median albedo contribution of 14.25%. Both the modeled and measured $D$ values decrease during solar maximum and increase toward solar minimum, consistent with the expected GCR modulation. A generally closer match is obtained for CRaTER before 2019, while noticeable discrepancies appear afterward, likely because the BON20 calibration precedes this period and 2019–2021 appears to be at an unusually quiet solar minimum. For TGO/Liulin-MO, the modeled (orange) and measured (brown) absorbed dose rates at Mars’s orbit show overall aligned temporal variations. In the modeled results, in addition to the correction factors for the presence of ACRs and the radial gradients of GCRs and ACRs in A, albedo particle contributions are accounted for using the fractions presented in Section 3.3 of Liu et al. (2023), which vary in the range of 4–8% depending on the solar modulation at an altitude of 400 km around Mars. As the modeled albedo particle contributions in Liu et al. (2023) provide specific values, the associated uncertainties are not shown here. However, additional uncertainties may arise from the variability of the model inputs (e.g., SSN) and from the regolith composition and physical properties of the Martian surface (e.g., Röstel et al., 2020), which influence the generation and transport of albedo particles and is beyond the scope of this study. The measured data are converted to dose rates in water with a factor of 1.333, as derived in Appendix B of Schwadron et al. (2012). Both modeled and measured values for TGO/Liulin-MO are systematically higher than those for CRaTER, due to (1) enhanced contributions from secondary particles under the thicker shielding environment (see also, e.g., Dobynde et al., 2019; Xu et al., 2022; Zhang et al., 2022; Liu et al., 2023), and (2) radial gradient of the GCR and ACR fluxes between Earth and Mars, as detailed in A.
@@ 4.1 Radiation Dose Contribution from Individual GCR Elements
In summary, the modulation of $\langle Q\rangle_{Z}$ with solar activity depends strongly on elemental mass and charge. In order to elucidate the underlying physics of the contrasting behavior mentioned above, we analyze how $\langle Q\rangle_{Z}$ evolves with solar modulation for light and heavy species in the following.
@@ Acknowledgments
The authors acknowledge the support by the National Natural Science Foundation of China (Grant Nos. 42521007, 42188101, 42474221, and W2433101). The authors thank Cary Zeitlin for his helpful suggestions on the CRaTER dataset and also express their gratitude to the groups developing the BON20 GCR model and the GEANT4 toolkit, the World Data Center SILSO, Royal Observatory of Belgium, Brussels, and the teams of ACE/CRIS, AMS-02, LRO/CRaTER and TGO/Liulin-MO instruments for providing valuable data involved in this study.
@@ Appendix A Calculations of ACR Contributions, and GCR and ACR Radial Gradients
Note that BON20 only incorporates GCRs, whereas ${D}_{\text{BON20}}^{*}(t)$ in Equation (3) includes both ACR and GCR contributed doses:
@@
and the dose equivalent $H$ is corrected in the same manner. For convenience, we compute the time-averaged correction factors for the GCR and ACR contributions in Equation (13) and apply these averaged factors to absorbed dose and dose equivalent estimates directly derived from BON20 in deep space or under different shielding conditions:
$\begin{aligned} &C_{r,\,{\rm GCR+ACR}}=C_{r,\,{\rm GCR}}+C_{r,\,{\rm ACR}}\\ &=\left\langle\frac{D_{\rm GCR}(r,t)}{D_{\rm BON20}(t)}\right\rangle_{t}+\left\langle\frac{D_{\rm ACR}(r,t)}{D_{\rm BON20}(t)}\right\rangle_{t}.\end{aligned}$ (14)
At $r=1.5$ AU ($\Delta r=0.5$ AU), the correction yields a 1.5% increase in both $D$ and $H$ due to GCR gradients, denoted as $C_{r,\,{\rm GCR}}$, and ($9.5\pm 0.06$)% and ($8.2\pm 1.3$)% increases in $D$ and $H$, respectively, due to ACR and its gradients, denoted as $C_{r,\,{\rm ACR}}$. Interestingly, at 1 AU, Equation (14) gives ($5.2\pm 0.06$)% and ($5.7\pm 0.7$)% increases in $D$ and $H$, respectively, consistent with the 6.3% estimate from Equation (5), confirming that including only He, N, and O provides a reasonable approximation of total ACR effects. Therefore, we adopt correction factors of 11.0% for $D$ and 9.7% for $H$ for long-term deep-space and shielded dose rates at Mars, although some deviations may occur for different shielding thicknesses.